Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
mantovani, spinicelli, bersanetti, nardecchia, pinto, gouaty, derossi - 8:19 Friday 04 August 2023 (61112) Print this report
ITF recovery

Yesterday morning we restarted the ITF recovery.

We have tuned the DARM phase for B1p 56MHz and the B2 169MHz phase along the lock acquisition.

We unfortunately unlocked all the time at 400mW (step 3) for an oscillation at 0.8 Hz for DARM, but if we increased the DARM gain we experienced high frequency oscillations.

Ilaria then changed very sligthly the CH power of 5% and we managed to survive 400mW and to tune the DIFFp demodulation phases (the change is very small and it is hard to explain the effect).

We managed then to have a Carm 0 3f (also the jump threshold has been changed).

Romain has also reduced the B1p QD1 galvo corrections with picomotors since the galvo was saturating (the one used for BS alignment).

to be continued today

Comments to this report:
derossi, mantovani, mwas, nardecchia, pinto, spinicelli - 18:23 Friday 04 August 2023 (61124) Print this report

This morning we continued the recovery of the lock acquisition after reaching CARM_NULL_3F yesterday night.

As a first step, we rechecked the calibration between B1p_PD1 and PD2, and we put the correct value (+0.071). Moreover, we also updated the jump trigger to a lower value (-0.00038). 

After that, we could reach CARM_NULL_1F. At this point, Michal locked the OMC and started to work on its alignment. We reached again LN2 at 10.25 UTC.

During the first ~2h of this lock, Michal continued the work on the OMC alignment. After this work,  at 12.15UTC we started to work on the CMRF optimisation by moving the setpoint of BS_TY.  We reached the minimum of the CMRF with a quite large offset on the BS_TY (-0.0013) with ~25Mpc. In these working conditions, we manually tuned the gain of SRCL in order to put its UGF @5Hz and we switched off all the DRMI lines (see trends in fig. 1).  

We then injected noise on DARM in order to have an estimation of the DCP with the minimum of CMRF:

  • 13.14.10UTC for 5'; "DARM_noise_broad" AMPL: 20 (to be noted: during the injection, we disabled SRCL_SET servo)

First estimation of the DCP gives ~395Hz.

Then, we went back with the BS_TY offset to have a comparison of the DCP without BS offset, but unfortunately, we unlocked at 13.26.UTC. Looking at the data (see fig. 2) we notice that there is a strange sudden saturation of LSC_SRCL_CORR exactly at the second 36. To be checked if it could be a problem related to data losses or similar.

In the following lock acquisition tries, we reached CARM null 3f where we unlocked twice for DARM low gain, and once at CARM null 1f for SDB1 drifting away. The last unlock could be related to a problem on the action of reload config in LSC_ACL (see comment below). 

To be noted: since we are now saturating the correction of the MICH_SET servo we increased the clipping thresholds in ACL from +/-10 to +/-20. 

 

Images attached to this comment
carbognani, poulton, bersanetti, spinicelli, pinto, sposito - 18:33 Friday 04 August 2023 (61126) Print this report

After changing the MICH_SET threshold in Acl, we clicked on 'reload config' on LSC_Acl process. This action seemed to create some trouble in the automation. After the unlock, indeed, while trying to restart the aquisition Metatron started to give some error messages of wrong selected state between some managed nodes (NARM/WARM_LOCK and ARMS_LOCK). After some investigation we found out that there was a flag that after the unlock didn't go to zero, that is LSC_ALS_NArm/WArm_SW being stucked at 1 (See Fig.1).

We tried to manually put the flag at 0 but the action didn't go through. Apparently, the action did before of reloading the config file in LSC_Acl process stucked some software trigger which stayed on also after the unlock.

By stopping and restarting LSC_Acl we were able to put manually the flag LSC_ALS_NArm/WArm_SW  = 0 and restart the acquisition.

Images attached to this comment
sorrentino - 13:09 Monday 07 August 2023 (61150) Print this report

Switching off DRMI lines reveals the noise background in the region between 20 Hz and 30 Hz, which is currently about a factor 2.5 above the O3 sensitivity

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×